### v4

# A Linear Representation for Functions on Finite Sets

Abstract
We demonstrate that any function $f$ from a finite set $Y$ to itself can be represented linearly. Specifically, we prove the existence of an injective map $j$ from $Y$ into a modular ring $\mathbb{Z}/m\mathbb{Z}$ and a constant $a\in\mathbb{Z}/m\mathbb{Z}$ such that $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ holds for all $y\in Y$. This result is established by analyzing the algebraic properties of the adjugate of the characteristic matrix associated with the function’s digraph. The proof is constructive, providing a method for finding the embedding $j$, the modulus $m$, and the linear multiplier $a$.
Keywords: Functional graph, Linear representation, Adjugate matrix, Characteristic polynomial, Finite functions.
1 Introduction
Functions on finite sets are fundamental objects in discrete mathematics, combinatorics, and theoretical computer science. The structure of such a function $f:Y\to Y$ is completely described by its functional graph (or functional digraph), a directed graph where a unique edge emanates from each vertex $y$ to its image $f(y)$. While these functions can exhibit complex and seemingly chaotic behavior, a central goal in mathematics is to find alternative representations that reveal underlying algebraic structure.
This article establishes that any such function becomes linear after a suitable embedding. We show that for any function $f$ on a finite set, one can find an injective map $j$ from the set into a ring of integers modulo $m$ such that the action of $f$ corresponds to multiplication by a constant $a$. That is for all $y\in Y$: $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ and hence the following diagram commutes:
[…]
This result bridges the gap between arbitrary discrete maps and structured linear systems. Without loss of generality, we can identify elements of the finite set $Y$ of size $n$ with the set $\{0,1,\ldots,n-1\}$, which are elements of $\mathbb{Z}/n\mathbb{Z}$. The full formal proof of this result is in [5].
2 Main Results
Definition 2.1 (Adjacency Matrix of a Function).
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. The function $f$ is represented by an $n\times n$ adjacency matrix $A=A_{f}$, where the entry $a_{ij}=\delta_{f(i),j}$ and $\delta_{i,j}$ is the Kronecker delta. With this convention, each row of AA contains exactly one non-zero entry.
Lemma $2.$
2.1Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Then for all $i\in\{0,1,\ldots,n-1\}$ and $y\in\mathbb{Z}^{n}$
$(Ay)_{i}=y_{f(i)}.$
Proof.
The proof follows from the Definition 2.1.
$(Ay)_{i}=\sum_{j=0}^{n-1}a_{ij}y_{j}=\sum_{j=0}^{n-1}\delta_{f(i),j}y_{j}=y_{f(i)}$
Lemma 2.3.
2.12.2Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Let $v\in\mathbb{Z}^{n}$, $y=\operatorname{adj}(xI-A)v$ and $m=\det(xI-A)$. Then for all $i\in\{0,1,\ldots,n-1\}$
$y_{f(i)}=xy_{i}-mv_{i}.$
Proof.
For adjugate matrix we have identity $(xI-A)\operatorname{adj}(xI-A)=\det(xI-A)I$. Therefore,
$mv=\det(xI-A)v=(xI-A)\operatorname{adj}(xI-A)v=(xI-A)y=xy-Ay.$
The final equality follows from the Lemma $2.$ ∎
Lemma 2.4.
Let $M$ be an $n\times n$ matrix with polynomial entries $m_{ij}\in\mathbb{Z}[x]$. Then
$\deg(\det(M))\leq\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}\deg(m_{ij}).$
Proof.
The determinant is a sum over permutations $\sigma$ of products $\prod_{i=0}^{n-1}m_{\sigma(i),i}$. Each product has degree at most $\sum_{i=0}^{n-1}\deg(m_{\sigma(i),i})$. Since a single element of a sum is at most the whole sum (when all terms are non-negative), this is bounded by $\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}\deg(m_{ij})$. The degree of a sum is at most the maximum degree of its summands. ∎
Lemma 2.5.
Let AA be an $n\times n$ matrix with integer entries. The characteristic matrix $\chi_{A}(x)=xI-A$ has determinant equal to the characteristic polynomial:
$\det(\chi_{A}(x))=\det(xI-A)$
For all $n\in N$, this polynomial is monic of degree $n$.
Proof.
This follows from the standard properties of the characteristic polynomial. ∎
Lemma 2.6.
2.4Let AA be an $n\times n$ matrix with integer entries and $\chi_{A}(x)=xI-A$ be its characteristic matrix. For $i\neq j$, the $(i,j)$ entry of $\operatorname{adj}(\chi_{A}(x))$ has degree at most n2n-2.
Proof.
The adjugate entry $\operatorname{adj}(\chi_{A}(x))_{ij}$ equals the determinant of the characteristic matrix with row $j$ and column $i$ removed from $\chi_{A}(x)$.
This submatrix has diagonal entries from $\chi_{A}(x)$ except at diagonal positions $i$ and $j$ (which are deleted). Since $\chi_{A}(x)$ has diagonal entries of degree 11 (from xIxI) and off-diagonal entries of degree 0 (from A-A), the submatrix has exactly n2n-2 diagonal entries of degree 11 and all other entries of degree 0.
By Lemma 2.4, the determinant has degree at most n2n-2. ∎
Lemma 2.7.
2.62.5Let $M=M(x)=\operatorname{adj}(xI-A)$ be the adjugate of the characteristic matrix xIAxI-A. Then the matrix entries $m_{ij}=p_{ij}(x)$ are polynomials in $x$ for all $i,j\in\{0,1,\ldots,n-1\}$ such that
  • $p_{ii}(x)$ is monic of degree n1n-1 for all $i\in\{0,1,\ldots,n-1\}$ and
  • $p_{ij}(x)$ has degree at most n2n-2 for all $i\neq j\in\{0,1,\ldots,n-1\}$.
Proof.
The diagonal entries of $\operatorname{adj}(xI-A)$ are characteristic polynomials of $(n-1)\times(n-1)$ submatrices, hence monic of degree n1n-1 by Lemma 2.5.
The off-diagonal case follows directly from Lemma 2.6. ∎
Definition 2.8.
For a polynomial $p(x)=\sum_{i=0}^{d}p_{i}x^{i}\in\mathbb{Z}[x]$, we define the coefficients bound:
$|p|=\sum_{i=0}^{d}|p_{i}|$
Lemma 2.9.
2.8If $p\in\mathbb{Z}[x]$ has positive leading coefficient, then for all integers $n\geq|p|$, we have $n>0$ and $p(n)>0$.
Proof.
Lemma is trivially true for $d=0$ so we can assume $d\geq 1$. Write $p(n)=an^{d}+r(n)$ where $a\geq 1$ is the leading coefficient of $p$, $d=\deg(p)$, and $\deg(r)<d$. For $n\geq|p|$:
$n\geq|p|\geq a\geq 1>0.$
Since $n\geq 1$, we have $|r(n)|\leq Bn^{d-1}$ where B=|p|aB=|p|-a. Therefore,
$p(n)=an^{d}+r(n)\geq an^{d}-Bn^{d-1}=n^{d-1}(an-B)\geq(a|p|-B)\geq|p|-B=a\geq 1>0.$
Lemma 2.10.
2.72.92.5Let $M=(m_{ij})=M(x)=\operatorname{adj}(xI-A)$ be the adjugate of the characteristic matrix xIAxI-A. Let $v=(1,2,\ldots,n)^{T}$ and $m=m(x)=\det(xI-A)$. Then for sufficiently large integer $x$:
$0<y_{0}<y_{1}<\cdots<y_{n-1}<m$
Proof.
The proof follows from Lemma 2.7 and Lemma 2.9. Let $y=Mv$. Then $y_{i}=\sum_{k=0}^{n-1}m_{ik}(k+1)$.
For each entry $y_{i}$, we express it as evaluation of a polynomial $p_{i}(x)=\sum_{k=0}^{n-1}m_{ik}(x)(k+1)\in\mathbb{Z}[x]$. By Lemma 2.7, the diagonal entry $m_{ii}$ is monic of degree n1n-1, while off-diagonal entries $m_{ik}$ (for $k\neq i$) have degree at most n2n-2. Therefore, the coefficient of $x^{n-1}$ in $p_{i}$ is $i+1>0$ (dominated by the $m_{ii}(i+1)$ term).
For the difference $p_{j}-p_{i}$ with $j>i$, the leading term comes from $(m_{jj}(j+1)-m_{ii}(i+1))$. Since both $m_{jj}$ and $m_{ii}$ are monic of degree n1n-1, the leading coefficient of $p_{j}-p_{i}$ is (j+1)(i+1)=ji>0(j+1)-(i+1)=j-i>0.
Similarly, for $p_{m}(x)=\det(xI-A)-p_{i}(x)$, since $\det(xI-A)$ is monic of degree $n$ (by Lemma 2.5) and $p_{i}$ has degree at most n1n-1, the leading coefficient is 1.
Since $p_{0}(x)$ has leading coefficient $0+1=1>0$, we have $y_{0}>0$ for sufficiently large $x$.
Applying Lemma 2.9 to these polynomials with positive leading coefficients gives the existence of $x_{0}$ such that all required inequalities hold for $x>x_{0}$. ∎
Definition 2.11 (Linear Representation).
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. A linear representation of $f$ is an injective function $j:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/m\mathbb{Z}$ such that for all $i\in\{0,1,\ldots,n-1\}$,
$j(f(i))=a\cdot j(i)$
in $\mathbb{Z}/m\mathbb{Z}$, where $m$ is a positive integer and $a$ is a constant from $\mathbb{Z}/m\mathbb{Z}$ depending on $f$.
Theorem 2.12 (Main Theorem).
2.112.32.10Any finite function $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ has a linear representation.
Proof.
Let us define $m=\det(xI-A)$, $a=x$ and $j(i)=y_{i}$ for $i\in\{0,1,\ldots,n-1\}$, where $y=Mv$ with $v=(1,2,\ldots,n)^{T}$. By Lemma 2.10, we can choose $x$ large enough such that:
$0<y_{0}<y_{1}<\cdots<y_{n-1}<m$
Since $m$ is monic of degree $n$, we can choose $x$ such that $x<m(x)$. This shows that $j$ is injective and $a\in\{0,1,\ldots,m-1\}$.
Lemma 2.3 shows that $y_{f(i)}=xy_{i}-mv_{i}$ and hence $y_{f(i)}\equiv xy_{i}\pmod{m}$.
Therefore, $j$ is a linear representation of $f$ with modulus $m$ and constant $a$. ∎
Examples
Example 2.13 (Quadratic Function in $\mathbb{Z}/3\mathbb{Z}$).
Consider the function $f:\mathbb{Z}/3\mathbb{Z}\to\mathbb{Z}/3\mathbb{Z}$ defined by $f(x)=x^{2}$. This function maps:
$0$ $\mapsto 0$
$1$ $\mapsto 1$
$2$ $\mapsto 4\equiv 1\pmod{3}$
Despite being a non-linear function, Theorem 2.12 guarantees that $f$ has a linear representation.
The adjacency matrix is:
$A_{f}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&1&0\end{pmatrix}$
The characteristic matrix is:
$xI-A_{f}=\begin{pmatrix}x-1&0&0\\ 0&x-1&0\\ 0&-1&x\end{pmatrix}$
The characteristic polynomial is:
$m=\det(xI-A_{f})=(x-1)^{2}\cdot x=x^{3}-2x^{2}+x$
The adjugate matrix is:
$\operatorname{adj}(xI-A_{f})=\begin{pmatrix}x(x-1)&0&0\\ 0&x(x-1)&0\\ 0&(x-1)&(x-1)^{2}\end{pmatrix}$
Using vector $v=(1,2,3)^{T}$, we get:
$y=\operatorname{adj}(xI-A_{f})\cdot v=\begin{pmatrix}x(x-1)\cdot 1\\ x(x-1)\cdot 2\\ (x-1)\cdot 2+(x-1)^{2}\cdot 3\end{pmatrix}=\begin{pmatrix}x^{2}-x\\ 2x^{2}-2x\\ (x-1)(3x-1)\end{pmatrix}=\begin{pmatrix}x^{2}-x\\ 2x^{2}-2x\\ 3x^{2}-4x+1\end{pmatrix}$
For $x=4$, we compute:
$y_{0}$ $=4^{2}-4=16-4=12$
$y_{1}$ $=2(4^{2})-2(4)=32-8=24$
$y_{2}$ $=3(4^{2})-4(4)+1=48-16+1=33$
$m$ $=4^{3}-2(4^{2})+4=64-32+4=36$
The injection $j:\mathbb{Z}/3\mathbb{Z}\to\mathbb{Z}/36\mathbb{Z}$ is defined by $j(i)=y_{i}$:
$j(0)=12,\quad j(1)=24,\quad j(2)=33$
These values are strictly increasing and bounded by $m=36$, so $j$ is injective.
We verify the linear representation property using Lemma 2.3.
The lemma states that $y_{f(i)}=xy_{i}-m\cdot v_{i}$, which we can rewrite as:
$j(f(i))\equiv xj(i)\pmod{m}$.
Verification:
$j(f(0))=j(0)=12$ $\equiv 4\cdot 12-36\cdot 1=48-36=12=4\cdot j(0)\pmod{36}\quad\checkmark$
$j(f(1))=j(1)=24$ $\equiv 4\cdot 24-36\cdot 2=96-72=24=4\cdot j(1)\pmod{36}\quad\checkmark$
$j(f(2))=j(1)=24$ $\equiv 4\cdot 33-36\cdot 3=132-108=24=4\cdot j(2)\pmod{36}\quad\checkmark$
Thus $j(f(i))\equiv 4\cdot j(i)\pmod{36}$ for all $i\in\mathbb{Z}/3\mathbb{Z}$, confirming the quadratic function $f(x)=x^{2}$ has a linear representation with modulus $m=36$ and multiplier $a=4$.
3 Discussion and Context
The representation of arbitrary functions in structured algebraic forms is a recurring theme. Our result provides one such representation, which can be compared to others.
  • Polynomial Interpolation: Any function over a finite field or ring can be represented by a polynomial (e.g., via Lagrange interpolation). Our result is distinct in that it achieves a representation as a simple linear map, $z\mapsto az$, at the cost of requiring an embedding $j$ into the ring rather than working on the original set’s labels directly.
  • Linear Feedback Shift Registers (LFSRs): In engineering and cryptography, functions on vector spaces over finite fields are often chosen to be linear for ease of implementation and analysis. Our result generalizes this principle, showing that any function on an arbitrary finite set can be viewed linearly in a suitable module.
  • Graph Theory: In Graph Theory our results shows that any functional digraph is a subgraph of a functional graph of a linear map.
While the modulus $m$ can be large, making this construction primarily of theoretical interest, it provides the powerful knowledge that every finite function possesses a hidden linear structure.